solve for x and show all the steps:
7(x+4)-10=3x+20+4x-2
Then, show all the steps to CHECK YOUR SOLUTION!
You will find something "odd" here that has to do with "infinity".
What is it?
What does it mean when "x" can be an "infinite" number of values?

Answer by robertb(5830)

7(x+4)-10=3x+20+4x-2
7x+28-10 = 7x+18
7x + 18 = 7x+18, ultimately
0 = 0.
What does this mean, my dear STUDENT? Obviously 0 = 0 is an irrefutably true statement (unless your number system is different from ours, hehe.) It means whatever real value you put in for x will always lead you to 0 = 0, which means the solution set is the set of ALL real numbers, which, you know, has infinitely many elements.